# Robust Kernel Hypothesis Testing under Data Corruption

### v1

@@ 3 Two-sample kernel testing robust to data corruption
Robust two-sample testing. We now consider the two-sample testing framework which, in the non-corrupted setting, satisfies the exchangeability assumption under the null [e.g., 50, Proposition 1]. In our robust setting, we are given samples $Y_{1},\ldots,Y_{n}$ and $Z_{1},\ldots,Z_{m}$, where at most $r$ of the $m+n$ samples have been corrupted (we do not know which ones, if any). Of the remaining samples (at least $m+n-r$ of them), the $Y_{i}$’s are drawn from $P$ while the $Z_{i}$’s are drawn from $Q$. We are interested in testing whether $P=Q$. The data corruption could potentially be applied in an adversarial way; for example, some of the $Y_{i}$’s samples could be replaced by new samples from $Q$. As before, the maximum number of corrupted samples $r$ is specified by the user and is considered to be known. We restrict ourselves to the case $r\leq n=\min(m,n)$, as the setting $r>n$ suffers from the same issue as when $r=n$: all the samples from $P$ could be corrupted, in which case all information about $P$ would be lost and we could not test whether $P=Q$.
@@
MMD statistic. As a divergence for this two-sample problem, we use the kernel-based MMD introduced by [29] which, for a kernel $k$, is defined as
$\mathrm{MMD}_{k}(P,Q)\coloneqq\sqrt{\mathbb{E}_{P}[k(Y,Y^{\prime})]-2\mathbb{E}_{P,Q}[k(Y,Z)]+\mathbb{E}_{Q}[k(Z,Z^{\prime})]}$
@@
for probability distributions $P$ and $Q$. The MMD is well-suited for the two-sample problem as it can distinguish between any two distributions in the sense that $\mathrm{MMD}_{k}(P,Q)=0$ if and only if $P=Q$, given that the kernel $k$ is characteristic [26]. For a given sample ${\mathcal{X}_{n+m}}\coloneqq(Y_{1},\ldots,Y_{n},Z_{1},\ldots,Z_{m})$ generated as described above, the quadratic-time empirical MMD statistic is
$\widehat{\mathrm{MMD}}(\mathcal{X}_{n+m})\coloneqq\bigg{(}\frac{1}{n^{2}}\sum_{i,j=1}^{n}k(Y_{i},Y_{j})+\frac{1}{m^{2}}\sum_{i,j=1}^{m}k(Z_{i},Z_{j})-\frac{2}{nm}\sum_{i=1}^{n}\sum_{j=1}^{m}k(Y_{i},Z_{j})\bigg{)}^{1/2},$
@@
which is a plug-in estimator, and its square can be expressed as a two-sample V-statistic. For a kernel bounded everywhere by $K$, [38, Lemma 5] provide an upper bound on the global sensitivity of this MMD statistic $\Delta_{\widehat{\mathrm{MMD}}}\leq\sqrt{2K}/\min(n,m)$, which is guaranteed to be tight when the kernel $k$ is translation invariant.
@@
dpMMD & dcMMD. We construct our robust two-sample dpMMD and dcMMD tests by applying the procedures of Algorithms 1 and 2 with the MMD statistic $T=\widehat{\mathrm{MMD}}$ and global sensitivity $\Delta_{T}=\sqrt{2K}/\min(n,m)$. It is immediately clear from Lemmas 1 and 2 that these robust tests control the type I error at the desired level $\alpha$ non-asymptotically, even under $r$ data corruption. We next prove the consistency of dpMMD and dcMMD in the data corruption framework, which ensures that any fixed distributions $P$ and $Q$ with $P\neq Q$ can be distinguished with probability one by these tests for large enough sample sizes. These tests scale quadratically with the sample sizes.
Lemma 3 (Consistency of dpMMD & dcMMD).
@@
Suppose that the kernel $k$ is characteristic, non-negative, and bounded everywhere by $K$. Assume that $n\leq m$, $r\varepsilon\leq\nu$ for some positive constant $\nu$ and $\max\{r/n,1/(\varepsilon n)\}\to 0$ as $n\to\infty$, and that a sequence of permutation numbers $B_{n}$ satisfies $\min_{n\in\mathbb{N}}B_{n}>c\alpha^{-1}-1$ with $c=e^{\nu}$ for dpMMD and with $c=1$ for dcMMD. Then, for any fixed $P$ and $Q$ with $P\neq Q$, both dpMMD and dcMMD are consistent in the sense that
@@
The consistency proof of Lemma 3 relies on proving that the sufficient conditions of Lemmas 1 and 2 are satisfied. We show that these conditions are met for the proposed MMD tests with characteristic and bounded kernels. Having obtained guarantees for the asymptotic power of dpMMD and dcMMD against fixed alternatives, we now consider the more challenging task of providing power guarantees that hold uniformly over classes of alternatives shrinking towards the null as the sample sizes increase. We choose these classes to be separated from the null in the MMD metric, and prove that dpMMD and dcMMD achieve the optimal uniform separation rate (with matching upper and lower bounds). This minimax result holds with respect to the smallest sample size, to the data corruption level, and to the quantities controlling the type I and II errors; hence, demonstrating the optimality of dcMMD and dpMMD in the data corruption setting.
Theorem 1 (Minimax optimal uniform separation of dpMMD & dcMMD).
Suppose that the kernel $k$ is characteristic, non-negative, and bounded everywhere by $K$. For $\alpha,\beta\in(0,1)$, assume that the number of permutations $B$ fulfills the condition in 3 in the appendix, and $n\asymp m$. Setting $\varepsilon=r^{-1}\max\{\log(e/\alpha),\log(e/\beta)\}$ for dpMMD, we have the following guarantees.
@@
    (Uniform separation) Both the dpMMD and dcMMD tests are guaranteed to have high power, i.e., $\mathbb{P}_{P,Q}\bigl{(}\textrm{\emph{reject $\mathcal{H}_{0}$}}\,|\,r\textrm{\emph{ corrupted data}}\bigr{)}\geq 1-\beta$ for any distributions $P$ and $Q$ separated as
    $\mathrm{MMD}_{k}(P,Q)\geq C_{K}\max\biggl{\{}\sqrt{\frac{\max\{\log(e/\alpha),\log(e/\beta)\}}{n}},\ \frac{r}{n}\biggr{\}}$
@@
Theorem 1 demonstrates that both the dpMMD and dcMMD tests achieve minimax rate optimality in terms of the MMD metric. In addition to this optimality, the dpMMD test inherently guarantees differential privacy, which is a side benefit worth highlighting. Further remarks regarding Theorem 1 are in order.
Low corruption regime. We first note that when $r\lesssim\sqrt{n\max\{\log(e/\alpha),\log(e/\beta)\}}$, the dominating term is the first one, which is the minimax rate in the usual no-corruption setting [38, Theorem 8]. In this low corruption regime, our two robust tests, dpMMD and dcMMD, achieve the same minimax optimal rate as the standard MMD test in the non-corrupted setting. This demonstrates that we can obtain robustness against corruption of up to $r$ samples, where $r\lesssim\sqrt{n}$, without compromising the uniform separation rate in power.
High corruption regime. When $\sqrt{n\max\{\log(e/\alpha),\log(e/\beta)\}}\lesssim r<n$, our robust MMD tests achieve high power when $\mathrm{MMD}_{k}$ exceeds $r/n$, which is the optimal rate in this high corruption regime as we proved in Theorem 1(ii). It is worth pointing out that the separation rate $r/n$ in the high corruption regime is independent of $\alpha$ and $\beta$, unlike the one obtained in the low corruption regime. However, we emphasise that both $\alpha$ and $\beta$ cannot be taken to be arbitrarily small as this would break the condition $\sqrt{n\max\{\log(e/\alpha),\log(e/\beta)\}}\lesssim r$.
Total corruption regime. When $r=n$, the separation rate becomes an arbitrary constant. This makes the condition for $\mathrm{MMD}_{k}$ in Theorem 1(i) vacuous since $\mathrm{MMD}_{k}$ is bounded above by $\sqrt{2K}$ for any distributions $P$ and $Q$. In fact, in this regime, all samples from $P$ could be adversarially replaced by samples from $Q$, so there is no hope of distinguishing between the two distributions.
@@ 4 Independence kernel testing robust to data corruption
Robust independence testing. Another testing framework satisfying the exchangeability assumption under the null is the independence testing one [e.g., 4, Proposition 1]. The robust independence testing problem is defined as follows. Given paired samples $\mathcal{X}_{n}=\{(Y_{i},Z_{i})\}_{i=1}^{n}$ where at most $r\in[n]$ of them have been corrupted (no knowledge of which ones are corrupted, if any), and the remaining paired samples (at least $n-r$ of them) are drawn from some joint distribution $P_{Y\!Z}$. We test whether that joint distribution $P_{Y\!Z}$ is equal to the product of its marginals $P_{Y}\times P_{Z}$, that is, we test for independence among the non-corrupted paired samples. Again, the maximum number of corrupted samples $r$ is considered known, usually specified by the user.
@@
HSIC statistic. As a measure of dependence, we use the HSIC [31] which, for kernels $k$ and $\ell$, is defined as
$\mathrm{HSIC}_{k,\ell}(P_{Y\!Z})\!=\!\sqrt{\mathbb{E}_{P_{Y\!Z}}\!\big{[}k_{Y,Y^{\prime}}\ell_{Z,Z^{\prime}}\big{]}\!\!-\!2\mathbb{E}_{P_{Y\!Z}}\!\big{[}\mathbb{E}_{P_{Y}}[k_{Y,Y^{\prime}}]\mathbb{E}_{P_{Z}}[\ell_{Z,Z^{\prime}}]\big{]}\!\!+\!\mathbb{E}_{P_{Y}}\!\big{[}k_{Y,Y^{\prime}}\big{]}\mathbb{E}_{P_{Z}}\!\big{[}\ell_{Z,Z^{\prime}}\big{]}}$
@@
with the condensed notation $k_{Y,Y^{\prime}}$ and $\ell_{Z,Z^{\prime}}$ for $k(Y,Y^{\prime})$ and $\ell(Z,Z^{\prime})$. Provided that both kernels are characteristic [28], the HSIC characterises dependence in the sense that $\mathrm{HSIC}_{k,\ell}(P_{Y\!Z})=0$ if and only if $P_{Y\!Z}=P_{Y}\times P_{Z}$; hence, it is suitable as a building block for independence testing. Given paired samples $\mathcal{X}_{n}=\{(Y_{i},Z_{i})\}_{i=1}^{n}$ generated by the process described above, the empirical HSIC statistic is
$\widehat{\mathrm{HSIC}}(\mathcal{X}_{n})\leavevmode\nobreak\ =\leavevmode\nobreak\ \bigg{(}\frac{1}{n^{2}}\sum_{i,j=1}^{n}k_{i,j}\ell_{i,j}+\frac{1}{n^{4}}\sum_{i_{1},i_{2},j_{1},j_{2}=1}^{n}k_{i_{1},j_{1}}\ell_{i_{2},j_{2}}-\frac{2}{n^{3}}\sum_{i,j_{1},j_{2}=1}^{n}k_{i,j_{1}}\ell_{i,j_{2}}\bigg{)}^{1/2},$
@@
where $k_{i,j}$ and $\ell_{i,j}$ denote $k(Y_{i},Y_{j})$ and $\ell(Z_{i},Z_{j})$ for $i,j\in[n]$. This quadratic-time plug-in estimator can be written as a one-sample fourth-order V-statistic. For kernels $k$ and $\ell$ bounded everywhere by $K$ and $L$, respectively, the global sensitivity of the HSIC statistic is bounded as $\Delta_{\widehat{\mathrm{HSIC}}}\leq 4\sqrt{KL}(n-1)/n^{2}$, which is asymptotically tight as shown in [38, Lemma 6].
@@
dpHSIC & dcHSIC. Under the data corruption setting, the robust dpHSIC and dcHSIC are constructed by applying the procedures of Algorithms 1 and 2 with the HSIC statistic $T=\widehat{\mathrm{HSIC}}$ and global sensitivity $\Delta_{T}$ as $4\sqrt{KL}(n-1)/n^{2}$. The validity of both tests is guaranteed by Lemmas 1 and 2 with non-asymptotic type I error control at the desired level $\alpha$ under $r$ data corruption. The next lemma proves that the dpHSIC and dcHSIC tests are consistent against any fixed alternative $P_{Y\!Z}$ with $P_{Y\!Z}\neq P_{Y}\times P_{Z}$. In other words, these tests can detect any fixed dependence between $Y$ and $Z$ with probability one for large enough sample size. These tests scale quadratically with $n$.
Lemma 4 (Consistency of dpHSIC and dcHSIC).
@@
Suppose that the kernels $k$ and $\ell$ are characteristic, non-negative, and bounded everywhere by $K$ and $L$, respectively. Assume that $r\varepsilon\leq\nu$ for some positive constant $\nu$ and $\max\{r/n,1/(\varepsilon n)\}\to 0$ as $n\to\infty$, and that a sequence of permutation numbers $B_{n}$ satisfies $\min_{n\in\mathbb{N}}B_{n}>c\alpha^{-1}-1$ with $c=e^{\nu}$ for dpHSIC and with $c=1$ for dcHSIC. Then, for any fixed joint distribution $P_{Y\!Z}$ with $P_{Y\!Z}\neq P_{Y}\times P_{Z}$, both dpHSIC and dcHSIC are consistent in the sense that
@@
Similarly to the robust two-sample testing setting, we can also obtain power guarantees for dpHSIC and dcHSIC in terms of uniform separation rates. More precisely, the type II errors of both tests can be controlled non-asymptotically whenever $\mathrm{HSIC}_{k,\ell}(P_{Y\!Z})$ is larger than the minimax rate (which tends to zero as the sample size grows), highlighting the optimality of dpHSIC and dcHSIC under the data corruption framework.
Theorem 2 (Minimax optimal uniform separation of dpHSIC & dcHSIC).
Suppose that the kernels $k$ and $\ell$ are characteristic, non-negative, and bounded everywhere by $K$ and $L$, respectively. For $\alpha,\beta\in(0,1)$, assume that the number of permutations $B$ fulfills the condition in 3 in the appendix. Setting $\varepsilon=r^{-1}\max\{\log(e/\alpha),\log(e/\beta)\}$ for dpHSIC, we have the following guarantees.
@@
    (Uniform separation) Both the dpHSIC and dcHSIC tests are guaranteed to have high power, i.e., $\mathbb{P}_{P_{Y\!Z}}\bigl{(}\textrm{\emph{reject $\mathcal{H}_{0}$}}\,|\,r\textrm{\emph{ corrupted data}}\bigr{)}\geq 1-\beta$ for any joint distribution $P_{Y\!Z}$ separated as
    $\mathrm{HSIC}_{k,\ell}(P_{Y\!Z})\geq C_{K,L}\max\biggl{\{}\sqrt{\frac{\max\{\log(e/\alpha),\log(e/\beta)\}}{n}},\ \frac{r}{n}\biggr{\}}$
@@
The interpretation of the independence uniform separation rate of Theorem 2 mirrors the one following Theorem 1 for the two-sample case. In the low corruption regime with $r\lesssim\sqrt{n\max\{\log(e/\alpha),\log(e/\beta)\}}$, robustness comes for free as the minimax rate is the same as in the non-corrupted setting. In the high corruption regime with $\sqrt{n\max\{\log(e/\alpha),\log(e/\beta)\}}\lesssim r<n$, only the alternatives for which $\mathrm{HSIC}_{k,\ell}(P_{Y\!Z})$ exceeds $r/n$ can be detected with high power, and this rate is optimal with respect to both $n$ and $r$ by Theorem 2(ii). In the total corruption regime where $r=n$, the uniform separation condition is never satisfied and any test is non-informative as all information is lost. Finally, we highlight again that our dpHSIC and dcHSIC tests are minimax optimal across all regimes of data corruption.

### v3

@@ 3 TWO-SAMPLE KERNEL TESTING ROBUST TO DATA CORRUPTION
Robust two-sample testing. We now consider the two-sample testing framework which, in the non-corrupted setting, satisfies the exchangeability assumption under the null (e.g., Schrab et al.,, 2023, Proposition 1). In our robust setting, we are given samples $Y_{1},\ldots,Y_{n}$ and $Z_{1},\ldots,Z_{m}$, where at most $r$ of the $m+n$ samples have been corrupted (we do not know which ones, if any). Of the remaining samples (at least $m+n-r$ of them), the $Y_{i}$’s are drawn from $P$ while the $Z_{i}$’s are drawn from $Q$. We are interested in testing whether $P=Q$. The data corruption could potentially be applied in an adversarial way; for example, some of the $Y_{i}$’s samples could be replaced by new samples from $Q$. As before, the maximum number of corrupted samples $r$ is specified by the user and is considered to be known. We restrict ourselves to the case $r\leq n=\min(m,n)$, as the setting $r>n$ suffers from the same issue as when $r=n$: all the samples from $P$ could be corrupted, in which case all information about $P$ would be lost and we would not be able to test whether $P=Q$.
@@
MMD statistic. As a divergence for this two-sample problem, we use the kernel MMD introduced by Gretton et al., 2012a which, for a kernel $k$, is defined as
$\mathrm{MMD}_{k}(P,Q)$
[…]
@@
for probability distributions $P$ and $Q$. The MMD is well-suited for the two-sample problem as it can distinguish between any two distributions in the sense that $\mathrm{MMD}_{k}(P,Q)=0$ if and only if $P=Q$, given that the kernel $k$ is characteristic (Fukumizu et al.,, 2008). For a given sample ${\mathcal{X}_{n+m}}\coloneqq(Y_{1},\ldots,Y_{n},Z_{1},\ldots,Z_{m})$ generated as described above, the quadratic-time empirical MMD statistic is
$\widehat{\mathrm{MMD}}(\mathcal{X}_{n+m})$
[…]
$-\frac{2}{nm}\sum_{i=1}^{n}\sum_{j=1}^{m}k(Y_{i},Z_{j})\bigg{)}^{1/2}$
@@
which is a plug-in estimator, and its square can be expressed as a two-sample V-statistic. For a kernel bounded everywhere by $K$, Kim and Schrab, (2023, Lemma 5) provide an upper bound on the global sensitivity of this MMD statistic $\Delta_{\widehat{\mathrm{MMD}}}\leq\sqrt{2K}/\min(n,m)$, which is guaranteed to be tight when the kernel $k$ is translation invariant.
@@
dcMMD. We construct our robust two-sample dcMMD test by applying the procedure of Algorithm 1 with the MMD statistic $T=\widehat{\mathrm{MMD}}$ and global sensitivity $\Delta_{T}=\sqrt{2K}/\min(n,m)$. It is immediately clear from Lemma 1 that the robust dcMMD test controls the type I error at the desired level $\alpha$ non-asymptotically, even under $r$ data corruption. We next prove its consistency in the data corruption framework, which ensures that any fixed distributions $P$ and $Q$ with $P\neq Q$ can be distinguished with probability one by the dcMMD test for large enough sample sizes. This test scales quadratically with the sample sizes.
Lemma 2 (Consistency of dcMMD).
@@
Suppose that the kernel $k$ is characteristic, non-negative, and bounded everywhere by $K$. Assume that $n\leq m$ and $r/n\to 0$ as $n\to\infty$, and that a sequence of permutation numbers $B_{n}$ satisfies $\min_{n\in\mathbb{N}}B_{n}>\alpha^{-1}-1$. Then, for any fixed $P$ and $Q$ with $P\neq Q$, dcMMD is consistent in the sense that
@@
The proof of Lemma 2 relies on proving that the sufficient condition of Lemma 1 is satisfied. We show that this condition is met for the proposed dcMMD test with characteristic and bounded kernel. Having obtained guarantees for the asymptotic power of dcMMD against fixed alternatives, we now consider the more challenging task of providing power guarantees that hold uniformly over classes of alternatives shrinking towards the null as the sample sizes increase. We choose these classes to be separated from the null in the MMD metric, and prove that dcMMD achieves the optimal uniform separation rate (with matching upper and lower bounds). This minimax result holds with respect to the smallest sample size, to the data corruption level, and to the quantities controlling the type I and II errors; hence, demonstrating the optimality of dcMMD in the data corruption setting.
Theorem 1 (Minimax optimal uniform separation of dcMMD).
Suppose that the kernel $k$ is characteristic, non-negative, and bounded everywhere by $K$. For $\alpha,\beta\in(0,1)$, assume that the number of permutations $B$ is greater than $3\alpha^{-2}\bigl{\{}\log\bigl{(}8/\beta\bigr{)}+\alpha(1-\alpha)\}$, and that $n\asymp m$ with $n\leq m$.
@@
    (Uniform separation) The dcMMD test is guaranteed to have high power, i.e., $\mathbb{P}_{P,Q}\bigl{(}\textrm{\emph{reject $\mathcal{H}_{0}$}}\,|\,r\textrm{\emph{ corrupted data}}\bigr{)}\geq 1-\beta$ for any distributions $P$ and $Q$ separated as
    $\mathrm{MMD}_{k}(P,Q)$
[…]
@@
Theorem 1 demonstrates that the dcMMD test achieves minimax rate optimality in terms of the MMD metric. Further remarks regarding Theorem 1 are in order, which we present below.
Low corruption regime. We first note that when $r\lesssim\sqrt{n\max\{\log(e/\alpha),\log(e/\beta)\}}$, the dominating term in the condition of Theorem 1(i) is the first one, which is the minimax rate in the usual no-corruption setting (Kim and Schrab,, 2023, Theorem 8). In this low corruption regime, the two robust dcMMD tests achieve the same minimax optimal rate as the standard MMD test in the non-corrupted setting. This demonstrates that we can obtain robustness against corruption of up to $r$ samples, where $r\lesssim\sqrt{n\max\{\log(e/\alpha),\log(e/\beta)\}}$, without compromising power in terms of uniform separation rate.
High corruption regime. When we have $\sqrt{n\max\{\log(e/\alpha),\log(e/\beta)\}}\lesssim r<n$, our robust dcMMD test achieves high power when $\mathrm{MMD}_{k}$ exceeds $r/n$, which is the optimal rate in this high corruption regime as we have proved in Theorem 1(ii). It is worth pointing out that the separation rate $r/n$ in the high corruption regime is independent of $\alpha$ and $\beta$, unlike the one obtained in the low corruption regime. However, we emphasise that both $\alpha$ and $\beta$ cannot be taken to be arbitrarily small as this would break the condition $\sqrt{n\max\{\log(e/\alpha),\log(e/\beta)\}}\lesssim r$.
Total corruption regime. When $r=n$ (or $r\geq n$), the separation rate becomes an arbitrary constant. This makes the condition for $\mathrm{MMD}_{k}$ in Theorem 1(i) vacuous since $\mathrm{MMD}_{k}$ is bounded above by $\sqrt{2K}$ for any distributions $P$ and $Q$. In fact, in this regime, all samples from $P$ could be adversarially replaced by samples from $Q$, so there is no hope of distinguishing between the two distributions.
@@ 4 INDEPENDENCE KERNEL TESTING ROBUST TO DATA CORRUPTION
Robust independence testing. Another testing framework satisfying the exchangeability assumption under the null is the independence testing one (e.g., Albert et al.,, 2022, Proposition 1). The robust independence testing problem is defined as follows. Given paired samples $\mathcal{X}_{n}=\big{(}(Y_{i},Z_{i})\big{)}_{i=1}^{n}$ where at most $r\in[n]$ of them have been corrupted (no knowledge of which ones are corrupted, if any), and the remaining paired samples (at least $n-r$ of them) are drawn from some joint distribution $P_{Y\!Z}$. We test whether that joint distribution $P_{Y\!Z}$ is equal to the product of its marginals $P_{Y}\times P_{Z}$, that is, we test for independence among the non-corrupted paired samples. Again, the maximum number of corrupted samples $r$ is considered known, usually specified by the user.
@@
HSIC statistic. As a measure of dependence, we use the HSIC (Gretton et al., 2005b, ) which, for kernels $k$ and $\ell$, is defined as
$\mathrm{HSIC}_{k,\ell}(P_{Y\!Z})\coloneqq\biggl{(}$ $\mathbb{E}_{P_{Y\!Z}}\!\big{[}k_{Y,Y^{\prime}}\ell_{Z,Z^{\prime}}\big{]}$
$-2\mathbb{E}_{P_{Y\!Z}}\!\big{[}\mathbb{E}_{P_{Y}}[k_{Y,Y^{\prime}}]\mathbb{E}_{P_{Z}}[\ell_{Z,Z^{\prime}}]\big{]}$
$+\mathbb{E}_{P_{Y}}\!\big{[}k_{Y,Y^{\prime}}\big{]}\mathbb{E}_{P_{Z}}\!\big{[}\ell_{Z,Z^{\prime}}\big{]}\biggr{)}^{1/2}$
@@
with the condensed notation $k_{Y,Y^{\prime}}$ and $\ell_{Z,Z^{\prime}}$ for $k(Y,Y^{\prime})$ and $\ell(Z,Z^{\prime})$. Provided that both kernels are characteristic (Gretton,, 2015), the HSIC characterises dependence in the sense that $\mathrm{HSIC}_{k,\ell}(P_{Y\!Z})=0$ if and only if $P_{Y\!Z}=P_{Y}\times P_{Z}$; hence, it is suitable as a building block for independence testing. Given paired samples $\mathcal{X}_{n}=\big{(}(Y_{i},Z_{i})\big{)}_{i=1}^{n}$ generated by the process described above, the empirical HSIC statistic is
$\widehat{\mathrm{HSIC}}(\mathcal{X}_{n})~{}\coloneqq~{}\bigg{(}$ $\frac{1}{n^{2}}\sum_{i,j=1}^{n}k_{i,j}\ell_{i,j}-\frac{2}{n^{3}}\sum_{i,j_{1},j_{2}=1}^{n}k_{i,j_{1}}\ell_{i,j_{2}}$
$+\frac{1}{n^{4}}\sum_{i_{1},i_{2},j_{1},j_{2}=1}^{n}k_{i_{1},j_{1}}\ell_{i_{2},j_{2}}\bigg{)}^{1/2}$
@@
where $k_{i,j}$ and $\ell_{i,j}$ denote $k(Y_{i},Y_{j})$ and $\ell(Z_{i},Z_{j})$ for $i,j\in[n]$. This quadratic-time plug-in estimator can be written as a one-sample fourth-order V-statistic. For kernels $k$ and $\ell$ bounded everywhere by $K$ and $L$, respectively, the global sensitivity of the HSIC statistic is bounded as $\Delta_{\widehat{\mathrm{HSIC}}}\leq 4\sqrt{KL}(n-1)/n^{2}$, which is asymptotically tight as shown by Kim and Schrab, (2023, Lemma 6).
@@
dcHSIC. Under the data corruption setting, the robust dcHSIC is constructed by applying the DC procedure of Algorithm 1 with the HSIC statistic $T=\widehat{\mathrm{HSIC}}$ and global sensitivity $\Delta_{T}$ as $4\sqrt{KL}(n-1)/n^{2}$. The validity of dcHSIC is guaranteed by Lemma 1 with non-asymptotic type I error control at the desired level $\alpha$ under $r$ data corruption. The next lemma proves that the dcHSIC test is consistent against any fixed alternative $P_{Y\!Z}$ with $P_{Y\!Z}\neq P_{Y}\times P_{Z}$. In other words, this test can detect any fixed dependence between $Y$ and $Z$ with probability one for large enough sample size. The dcHSIC test scales quadratically with $n$.
Lemma 3 (Consistency of dcHSIC).
@@
Suppose that the kernels $k$ and $\ell$ are characteristic, non-negative, and bounded everywhere by $K$ and $L$, respectively. Assume that $r/n\to 0$ as $n\to\infty$, and that a sequence of permutation numbers $B_{n}$ satisfies $\min_{n\in\mathbb{N}}B_{n}>\alpha^{-1}-1$. Then, for any fixed joint distribution $P_{Y\!Z}$ with $P_{Y\!Z}\neq P_{Y}\times P_{Z}$, dcHSIC is consistent in the sense that
@@
Similarly to the robust two-sample testing setting, we can also obtain power guarantees for dcHSIC in terms of uniform separation rates. More precisely, the type II errors of dcHSIC can be controlled non-asymptotically whenever $\mathrm{HSIC}_{k,\ell}(P_{Y\!Z})$ is larger than the minimax rate (which tends to zero as the sample size grows), highlighting the optimality of dcHSIC under the data corruption framework.
Theorem 2 (Minimax optimal uniform separation of dcHSIC).
Suppose that the kernels $k$ and $\ell$ are characteristic, non-negative, and bounded everywhere by $K$ and $L$, respectively. For $\alpha,\beta\in(0,1)$, assume that the number of permutations $B$ is greater than $\alpha^{-2}\bigl{\{}\log\bigl{(}8/\beta\bigr{)}+\alpha(1-\alpha)\}$.
@@
    (Uniform separation) The dcHSIC test is guaranteed to have high power, i.e., $\mathbb{P}_{P_{Y\!Z}}\bigl{(}\textrm{\emph{reject $\mathcal{H}_{0}$}}\,|\,r\textrm{\emph{ corrupted data}}\bigr{)}\geq 1-\beta$ for any joint distribution $P_{Y\!Z}$ separated as
    $\mathrm{HSIC}_{k,\ell}(P_{Y\!Z})$
[…]
@@
The interpretation of the independence uniform separation rate of Theorem 2 mirrors the one following Theorem 1 for the two-sample case. In the low corruption regime with $r\lesssim\sqrt{n\max\{\log(e/\alpha),\log(e/\beta)\}}$, robustness comes for free as the minimax rate is the same as in the non-corrupted setting. In the high corruption regime with $\sqrt{n\max\{\log(e/\alpha),\log(e/\beta)\}}\lesssim r<n$, only the alternatives for which $\mathrm{HSIC}_{k,\ell}(P_{Y\!Z})$ exceeds $r/n$ can be detected with high power, and this rate is optimal with respect to both $n$ and $r$ by Theorem 2(ii). In the total corruption regime where $r=n$, the uniform separation condition is never satisfied and any test is non-informative as all information is lost. Finally, we highlight again that our dcHSIC test is minimax optimal across all regimes of data corruption.
5 RELATED WORK: ROBUSTNESS VIA DIFFERENTIAL PRIVACY
Robust tests can also be constructed from existing differentially private tests. In Section 5.1, we present background information on differential privacy. In Section 5.2, we construct robust tests by leveraging the existing differential privatisation (DP) method of Kim and Schrab, (2023) for permutation tests. We prove their validity and defer its power analysis to Appendix D.
5.1 Differential privacy background
Differential privacy (Dwork et al.,, 2006) is a popular framework for private data analysis that aims to return an output insensitive to any single individual’s data point. As discussed in Dwork and Lei, (2009); Avella-Medina, (2020), there is a close connection between robust statistics and differential privacy. Both frameworks require that algorithms remain stable under data perturbation. Recent research has delved into extending non-private tests to privacy settings and studying their asymptotic properties (Fienberg et al.,, 2011; Couch et al.,, 2019; Campbell et al.,, 2018; Rogers and Kifer,, 2017; Raj et al.,, 2020) and non-asymptotic properties (Cai et al.,, 2017; Aliakbarpour et al.,, 2018, 2019; Acharya et al.,, 2019; Kim and Schrab,, 2023).
Differential privacy (DP) is closely related to statistical robustness (Dwork and Lei,, 2009; Avella-Medina,, 2020). This connection naturally motivates us to leverage the recent advances in differential privatisation of hypothesis tests (Peña and Barrientos,, 2022; Kazan et al.,, 2023; Kim and Schrab,, 2023) for our data corruption setting. Consider any valid $\varepsilon$-DP hypothesis test with adjusted lower level $\alpha e^{-r\varepsilon}$ and pure DP parameter $\varepsilon>0$. Under the null, the DP group privacy property (Dwork et al.,, 2014, Theorem 2.2) ensures that
$\sup_{P_{0}\in\mathcal{P}_{0}}\mathbb{P}_{P_{0}}\bigl{(}\textrm{reject }\mathcal{H}_{0}\,\,|\,\,r\textrm{ corrupted data}\bigr{)}$ (1)
[…]
We choose to focus on the differential privatisation approach of Kim and Schrab, (2023) for permutation tests, as it benefits from stronger theoretical guarantees. In particular, their resulting DP tests are guaranteed to be valid at every sample size (Kim and Schrab,, 2023, Theorem 1), so type I error control under data corruption of Equation 1 holds non-asymptotically. This is actually crucial as the corruption and privacy parameters $r$ and $\varepsilon$ usually depend on the sample size. Consequently, the adjusted level $\alpha e^{-r\varepsilon}$ also depends on the sample size, which can modify the asymptotic behaviour of type I error control. We simply refer to this test construction based on Kim and Schrab, (2023) with the adjusted level as the DP procedure (see details in Algorithm 2).
The consistency of the DP test against data corruption can be guaranteed (i.e., for a fixed alternative distribution, the test power converges to one asymptotically) under minimal conditions. Such a result is presented in Appendix D, alongside with a power separation analysis of the two-sample dpMMD and independence dpHSIC tests, constructed from Algorithm 2.
Algorithm 2 Robust DP procedure
Adapted from Kim and Schrab,, 2023
Inputs: Data $\mathcal{X}_{n}$, robustness $r$, level $\alpha$, statistic
$T$, permutation number $B$, privacy $\varepsilon$.
Generate i.i.d. $\zeta_{0},\dots,\zeta_{B}\sim\mathsf{Laplace}(0,1)$.
Generate i.i.d. permutations $\bm{\pi}_{1},\dots,\bm{\pi}_{B}$ of $[n]$.
Set $\bm{\pi}_{0}=\mathrm{Id}$ and compute global sensitivity $\Delta_{T}$.
Compute $(1\!-\!\alpha e^{-r\varepsilon})$-quantile $q$ of $M_{0},\dots,M_{B}$.
Output: Reject $\mathcal{H}_{0}$ if $M_{0}>q$.